Water flows through a 4.70 cm diameter hose at 1.35×10−4 m3/s .

Part A

What's the flow speed?

Express your answer with the appropriate units.
Answer #1

flow rate = speed * Area

speed = flow rate / Area

area = pi*r^2

r = 4.70/2 cm = 2.35 cm = 2.35 x 10^-2 m

Area = pi * (2.35 x 10^-2)^2

speed = 1.35 x 10^- 4 / pi * (2.35 x 10^-2)^2

speed = 0.07785 m/s = 0.078 m/s